• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 08:01:13 on March28,2005

    Entry number 142631

    keyword=Shift Summary

    Started shift with run 3008

    All production runs at 40uA beam on C-tilted.

    00:26 Left Q1 tripped.

    00:28 Stopped the run due to Left Q1 tripp

    Remote reset of Q1 worked fine.
    00:51 Q1 back

    00:51-01:44 Run 3009(88k) , seen event rate only 25 Hz instead of ~75Hz. Stopped the run, run-end failed due to ROC1 problem. Rebooted ROC1. Did "startcoda", failed due to ROC1 again, ROC1 did not respond. Did reset ROC1. Now fine.

    01:54-02:26 Run 3010(88k), again repeated above problem. Did reset all ROCs and TSs including SRC ones too. Worked fine.

    02:38-03:23 Run 3011(200k).

    03:24-04:11 Run 3012(217k)
    04:11-05:08 Run 3013(243k)
    Run3014 junk, ROC6 failed responding, did reset.

    06:10-07:01 Run 3015(203k)
    07:02-07:50 Run 3016(226k)

    07:50- Run 3017 passed to day shift.